Differential Equations
Solve the given initial value problem. y - 2y - 36y + 72y = 0 y(O)= -13, y(O)= - 34y(0) = - 308 y(x) = 0
Answer #1

The given differential equation is y'''-2y''-36y'+72=0

with initial conditions y(0)=-13,\;y'(0)=-34,\;y''(0)=-308

This is a homogenous differential equation of 3rd order with constant coefficients. The auxiliary equation is

m^3-2m^2-36m+72=0\Rightarrow (m-2)(m^2-36)=0\\ \Rightarrow m=2,6,-6

Hence the general solution is

y(x)=c_1e^{2x}+c_2e^{-6x}+c_3e^{6x}

where c_1,c_2,c_3 are constants which can be determined using initial conditions.

If the general solution satisfies the initial condition, then we obtain the following equations

c_1+c_2+c_3=-13

2c_1-6c_2+6c_3=-34

4c_1+36c_2+36c_3=-308

We shall solve this linear system of equations by matrix row-echelon method with augmented matrix

\begin{pmatrix}1&1&1&-13\\ \:2&-6&6&-34\\ \:4&36&36&-308\end{pmatrix}

R_2\rightarrow R_2-2R_1, R_3\rightarrow R_3-4R_1

\begin{pmatrix}1&1&1&-13\\ \:0&-8&4&-8\\ \:0&32&32&-256\end{pmatrix}

R_2\rightarrow \frac{R_2}{4},\;R_3\rightarrow \frac{R_3}{32}

\begin{pmatrix}1&1&1&-13\\ \:0&-2&1&-2\\ \:0&1&1&-8\end{pmatrix}

R_3\rightarrow 2R_3+R_2

\begin{pmatrix}1&1&1&-13\\ \:\:0&-2&1&-2\\ \:\:0&0&3&-18\end{pmatrix}

This gives

c_1=-5,\;c_2=-2,\;c_3=-6

hence general solution is

y(x)=-5e^{2x}-2e^{-6x}-6e^{6x}